the perimeter of a rectangle is 110 meters. the width of the rectangle is 20 meters. how long is the rectangle?

Answer by farbonara(13)

A rectangle has 2 lengths and 2 widths.
Sum of 2 widths=20x2=40
Sum of 2 lengths=110-40=70
1 length=70÷2=35